OpenStudy (anonymous):

A square has an area of 16x^2+40x+25. Find the length of a side. Make a sketch of the square.

OpenStudy (anonymous):

area=x^2 16x^2+40x-25=x^2 15x^2+40x+25=0

OpenStudy (anonymous):

find x using quadratic formula or by factorization

OpenStudy (anonymous):

5(3x^2+8x+5)=0 5(x+1)(3x+5)=0 x=-1 or x=-3/5